OpenStudy (anonymous):

Ah! Please help! I'm about to lose power due to a very bad snow storm and this is the very last question I have from my online homework, and I just want to get this done before I am in the dark! A ball is thrown across a playing field from a height of h = 3 ft above the ground at an angle of 45° to the horizontal at the speed of 20 ft/s. It can be deduced from physical principles that the path of the ball is modeled by the function y= − (32 / (20)^2)x^2 + x + 3 (a) Find the maximum height attained by the ball. (Round your answer to three decimal places.)

OpenStudy (anonymous):

(b) Find the horizontal distance the ball has traveled when it hits the ground. (Round your answer to one decimal place.)

OpenStudy (surry99):

so this a projectile motion question ...agreed?

OpenStudy (anonymous):

Yes.

OpenStudy (surry99):

ok at it maximum height, what is the y component of the velocity of the ball?

OpenStudy (anonymous):

20 ft/s?

OpenStudy (surry99):

No, at the top the vertical component of the velocity must be equal to zero

OpenStudy (surry99):

you still there?

OpenStudy (anonymous):

Yes.

OpenStudy (surry99):

do you agree about the y component at the top?

OpenStudy (anonymous):

yes

OpenStudy (surry99):

great now the equation of motion describes an upside down parabola agreed?

OpenStudy (anonymous):

Yes.

OpenStudy (surry99):

ok, what is the slope of the parabola when the ball is at the top?

OpenStudy (surry99):

Draw a sketch if you need to...

OpenStudy (surry99):

still there

OpenStudy (anonymous):

Yes, sorry. I'm trying to figure it out but I seriously do not know.

OpenStudy (surry99):

That is ok...I will help you. The slope of the parabola woould be zero at the top. What does that mean in terms of dy/dx?

OpenStudy (surry99):

dy/dx would be equal to zero. Since you are given y = f(x) you could take the derivative and set it equal to zero to find the corresponding value of x....agreed?

OpenStudy (anonymous):

Yes.

OpenStudy (surry99):

ok go ahead and lets see what you get....

OpenStudy (anonymous):

I got 9.125 for the max height

OpenStudy (surry99):

so what value did you get for x and how did you get y...just want to make sure

OpenStudy (anonymous):

Used the formula k = c - (b^2/4a) *my professor said that is used to find the max height* but webassign wont take that answer. I mean, it's already rounded to three decimal places.

OpenStudy (surry99):

hang on

OpenStudy (surry99):

I have y = -(32/400)x^2 + x + 3 ...correct?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

I got it now! My professor just sent out a mass email saying he gave us the wrong equation to find the max height and to find the distance. Thank you for your help though.

OpenStudy (surry99):

so dy/dx = -.16x + 1 so 0 = -.16x +1 x=1/.16 = 6.25 then for the height y y = -.08(6.25^2) + 6.25 + 3 y = 6.125

OpenStudy (surry99):

you are ok then or do you want me to help you finish

OpenStudy (anonymous):

Nope, I got it. Thank you though. I'll give you a medal for helping me with this roller coaster of a question.

OpenStudy (surry99):

ok let me know if you need any help in the future. Take care and stay warm in the storm.

OpenStudy (anonymous):

Alright and will do!